let x=33/65 be the x-coordinate od the point P(x,y), where the terminal side of angle theta (in standard position) meets the unit circle. If P is in Quadrant IV, what is sin theta?

Answer by Alan3354(69443)

let x=33/65 be the x-coordinate od the point P(x,y), where the terminal side of angle theta (in standard position) meets the unit circle. If P is in Quadrant IV, what is sin theta?
---------------
x = 33/65
x^2 + y^2 = 1
Solve for y.
================
sin = y/r = y/1 = y
y and sine are negative in Q4